OpenStudy (anonymous):

precalculus question: how do you know that you've found all the values of x when asked to find all values of x satisfying (for example) 2(cosx)^2-cosx=0 on the interval (0, 2pi)

OpenStudy (campbell_st):

this is a quadratic equation that can be simplified. cos(x)(2cos(x) - 1)=0 so cos(x) = 0 and 1/2 now you with need to know about the cos(x) curve or the quadratic... cos is positive in the 1st and 4th quadrants... here is a rough sketch of the y=cos(x)  use an exact value triangle to find cos(x) = 1/2 1st quadrant is just the angle, 4th quadrant is 2pi - angle. hope it helps